# DualNILM: Energy Injection Identification enabled Disaggregation with Deep Multi-Task Learning

### v2

@@
Abstract
Non-Intrusive Load Monitoring (NILM) offers a cost-effective method to obtain fine-grained appliance-level energy consumption in smart homes and building applications. However, the increasing adoption of behind-the-meter (BTM) energy sources such as solar panels and battery storage poses new challenges for conventional NILM methods that rely solely on at-the-meter data. The energy injected from the BTM sources can obscure the power signatures of individual appliances, leading to a significant decrease in NILM performance. To address this challenge, we present DualNILM, a deep multi-task learning framework designed for the dual tasks of appliance state recognition and injected energy identification. Using a Transformer-based architecture that integrates sequence-to-point and sequence-to-sequence strategies, DualNILM effectively captures multiscale temporal dependencies in the aggregate power consumption patterns, allowing for accurate appliance state recognition and energy injection identification. Extensive evaluation on self-collected and synthesized datasets demonstrates that DualNILM maintains an excellent performance for dual tasks in NILM, much outperforming conventional methods. Our work underscores the framework’s potential for robust energy disaggregation in modern energy systems with renewable penetration. Synthetic photovoltaic augmented datasets with realistic injection simulation methodology will be open-sourced after review.
I Introduction
Non-Intrusive Load Monitoring (NILM) technique enables disaggregation of total household energy consumption into individual appliance usage patterns [1], offering a cost-effective solution for fine-grained energy monitoring. This capability is increasingly critical for smart grid applications, facilitating targeted energy-saving recommendations, demand response programs, and enhanced grid efficiency [2].
Conventional NILM methods operate on a key assumption: the aggregate power measurements from the utility meter represent the sum of purely consumptive loads. Then these methods leverage unique electrical signatures of appliances to infer their operating states (or power) within the total load [3, 4, 5]. However, the rapid adoption of behind-the-meter (BTM) renewable energy systems, particularly solar photovoltaics (PV), fundamentally disrupts this premise. The injection of energy from BTM sources creates bidirectional power flows that can dominate the aggregate power profile (especially during peak generation periods) [6, 5], severely obscuring the signatures of individual appliances and challenging the foundation of traditional NILM.
@@ I Introduction
The technical complexity of this challenge is significant. When PV generation exceeds instantaneous consumption, the disaggregation problem undergoes a fundamental mathematical transformation. Traditional NILM benefits from non-negativity constraints that bound the solution space. Each appliance contributes positively to the aggregate, creating a convex optimization landscape. In contrast, BTM generation shatters these constraints, expanding the feasible region to include arbitrary combinations of positive consumption and negative generation that yield identical net measurements. This transformation is further complicated by the electrical characteristics of modern grid-tied inverters, which operate at near-unity power factors to maximize generation efficiency. While optimal for energy production, this operational mode eliminates the reactive power signatures that traditionally help distinguish different load types. This critical disambiguation channel is lost just when it is needed most.
Figure 1 empirically demonstrates these theoretical challenges through controlled laboratory experiments. Our micro-inverter injection setup reveals how even modest energy generation fundamentally distorts the aggregate signal. During injection periods (shown in green), the expected additive structure of load aggregation breaks down. The aggregate signal (orange) deviates dramatically from the consumption-only baseline (blue), while the actual appliance states (bottom panels) become completely obscured. This controlled experiment confirms that the challenge transcends simple noise addition, revealing a fundamental incompatibility between conventional NILM architectures and the bidirectional power flows of modern energy systems.
@@
Systematic characterization of BTM challenges: We analyze how behind-the-meter PV injection transforms the NILM problem from a constrained optimization to an under-determined inverse problem, formally identifying key failure modes such as signal eclipse and temporal aliasing.
@@
Comprehensive empirical validation: We demonstrate DualNILM’s state-of-the-art performance across diverse operational regimes, from controlled laboratory experiments (500W micro-inverter) to synthetic residential-scale scenarios (2kW PV systems).
Open research infrastructure: We release the synthetic PV-augmented versions (with realistic weather data) of REDD and UK-DALE datasets and the corresponding benchmark suites to support future research in this critical area.
II Related work
@@
Non-Intrusive Load Monitoring (NILM) has been a subject of extensive research since its introduction by Hart [1]. Traditional NILM approaches rely on identifying unique appliance signatures in the aggregate power signal to disaggregate individual appliance usage [3]. Techniques such as Hidden Markov Models (HMMs) [7], factorial HMMs [8], and combinatorial optimization [9] have been widely used. With the advent of deep learning, neural network-based methods have significantly advanced NILM performance. Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) have been employed to capture temporal dependencies and extract features from power signals [10, 11, 12]. Sequence-to-point and sequence-to-sequence models have been proposed to improve disaggregation accuracy [11, 13]. Recently, researchers have also introduced advanced deep architectures, such as transformer-based models and graph neural networks, to further enhance disaggregation performance [14, 15]. These approaches leverage long-range temporal context and inter-appliance relationships to achieve higher accuracy on benchmark datasets while improving the generalization ability of NILM models.
@@
NILM methods can also be categorized according to their task-oriented objectives. Many approaches focus on event-based detection, where the key aim is to recognize appliance switching events (ON/OFF transitions) and label those events with the corresponding device [16, 17]. This category typically relies on detecting changes in aggregate power that exceed a certain threshold, followed by a classification stage to identify which appliance caused the event. Various improvements to event detection have been proposed, including hybrid schemes that combine primary event triggers with secondary verification algorithms to reduce false detections [18]. By isolating transition points in time series, event-based NILM methods can be highly interpretable and computationally efficient, but may struggle with appliances that exhibit gradual power variations or multiple internal states [19]. On the other hand, the steady-state-based approaches emphasize continuous monitoring of power consumption, estimating the exact load of each device over time. These NILM methods often employ more complex models (e.g., HMMs or deep neural networks) and can capture multi-level operating states of appliances [20].
NILM Methods amid RE Integration
Previous work addressing NILM with the integration of renewable energy is limited. In [21], the authors investigate methods for managing power injection systems in renewable setups, emphasizing the stability of the system. Similarly, Aygen et al. [22] consider a zero-sequence current injection technique to mitigate unbalanced loads in systems integrated with renewables, which, however, is not directly applicable to NILM. Only a few studies have attempted to explicitly disaggregate solar generation from the net metered signal. Noori et al. [23] recently examined solar disaggregation and highlighted that incorporating an external context, such as solar irradiance, can improve the detection of solar generation. In [24], the authors investigate solar energy disaggregation but rely on high-frequency data and additional features, which may not be practical in real-world deployments. Other studies have considered the use of additional sensors or prior knowledge of the energy injection [25], which limits their applicability.
@@
III Problem Definition
Traditional NILM approaches infer appliance-level power consumption from aggregated household measurements under the assumption that all loads are strictly non-negative [1, 9]. However, modern residences increasingly incorporate BTM renewable generation, where micro-inverters inject power back to the grid, fundamentally violating this core assumption.
III-A NILM with BTM Energy Injection
@@
where aggregates remaining loads and noise. While solving for is NP-hard [1], the non-negative structure historically enables acceptable performance through heuristic and learning-based methods [3, 4].
@@
III-B DNN-based NILM with Dual Tasks
Deep neural networks (DNNs) have shown promise in NILM by exploiting additional features or electrical signatures beyond active power [10, 11]. Let denote the multi-dimensional feature vector at time , including active power, reactive power, current harmonics, etc. A neural model processes these features as:
@@ III-B DNN-based NILM with Dual Tasks
This formulation enables the model to learn complementary representations: discrete state transitions for appliances and continuous power variations for injection. By explicitly modeling injection rather than treating it as negative consumption, the framework naturally accommodates bidirectional power flows while reducing optimization complexity. The dual-task NILM formulation is also well-suited for DNN-based NILM models that can leverage rich features, capture complex patterns and dependencies in the data, allowing accurate estimation of the appliance states and the energy injection [10, 11, 26].
@@ IV-B Model Architecture
where is the -th feature channel and is the corresponding encoded representation. Using separate CNN Encoders for each feature allows the model to learn feature-specific local patterns, enhancing its ability to capture the unique characteristics of different electrical measurements [10].
@@
The outputs of the CNN Encoders are concatenated along the feature dimension and fed into Transformer Encoders [27]. The Transformer Encoder employs multi-head attention mechanisms to capture long-range dependencies in the input sequence:
where is the output of the Transformer Encoder. This component allows DualNILM to model complex interactions between different appliances and the energy injection source over time, which is crucial for accurate disaggregation and injection estimation [11].
@@
The proposed architecture addresses the signal eclipse phenomenon directly through the above coupled design choices, ensuring both theoretical soundness and practical viability. First, the CNN-Transformer hybrid accommodates distinct temporal characteristics: appliance switching events exhibit sharp transients captured by convolutional kernels, while energy injection exhibits complex multiscale temporal dependencies and cross-correlation patterns requiring long-range attention mechanisms to capture their dynamics. Second, explicit separation of injection estimation through dedicated Seq2Seq projection prevents mathematical contamination of appliance state classification, where negative flows would corrupt binary decision boundaries. This architectural separation, combined with LayerNorm stabilization [28], enables direct processing of raw meter readings while maintaining modular scalability for incremental appliance addition. Third, the shared encoder representations enable cross-task synergy that significantly reduces the problem of identification, allowing accurate disambiguation between actual load reduction and injection-induced power dips within their respective constraint domains.
IV-C Training Objective and Optimization
@@
Our model can be optimized using gradient backpropagation approaches such as SGD [29], Adam [30], etc. Specifically, we employ Dice loss [31], which is particularly effective for imbalanced binary classification problems for appliance state recognition:
For injection estimation loss, we use the standard L2 measure: Root Mean Square Error (RMSE). Note that we can apply constraints/normalization on injection disaggregation targets as the maximum output of micro-inverters or community solar installations typically must adhere to standardized values like EN50583 [32].
V Experimental Setup
@@ V-A Datasets and Simulation
We collected a real-world dataset in our laboratory and conducted extensive experiments. The dataset includes individual power consumption data for multiple appliances and the BTM energy injection from a micro-inverter. The data was collected from December 6 to December 15, 2023, and was sampled at 2-second intervals using sub-meters for each appliance on both active power and reactive power. The dataset includes the following appliances: i) Air purifier, ii) Heater, iii) Light bulb Type I, iv) Light bulb Type II, v) Air compressor, vi) Air conditioner, vii) Micro-inverter (energy injection). Table I presents the key statistics of our laboratory dataset.
@@
To evaluate the generalizability of DualNILM and facilitate comparison with existing methods, we simulate energy injection scenarios using two widely-used public NILM datasets: REDD [9] and UK-DALE [33]. Our simulation takes advantage of real-world solar irradiance data and photovoltaic (PV) system modeling to create realistic scenarios of residential solar energy injection. Detailed implementation will be open-source and available after acceptance. The simulation process involved the following steps:
1.
@@
2.
Acquisition of Solar Irradiance Data: We obtain historical solar irradiance data from the National Solar Radiation Database (NSRDB) provided by the National Renewable Energy Laboratory (NREL) [34]. The NSRDB offers detailed solar radiation data, including Global Horizontal Irradiance (GHI), Direct Normal Irradiance (DNI), and Diffuse Horizontal Irradiance (DHI), with a temporal resolution of 30 minutes. We collect data for the rough geographical locations corresponding to the datasets:
@@
3.
Simulation of Photovoltaic System Output: Using the collected irradiance data, we simulate a residential PV system with a designed maximum output capacity of 120 W, representing a small-scale household solar installation. The simulation accounts for the effects of temperature on the efficiency of PV, following the methods described in [35, 36].
@@
(16) where is the rated capacity (120 W in our simulation) and is the inverter efficiency (0.96 based on [37]).
4.
Reactive Power Simulation: For comprehensive power analysis, we simulate reactive power characteristics for both appliances and the PV system based on typical power factors from industry standards [38]. For each appliance with active power , its reactive power is calculated as:

### v3

# Energy Injection Identification enabled Disaggregation with Deep Multi-Task Learning

@@
Abstract.
Non-Intrusive Load Monitoring (NILM) offers a cost-effective method to obtain fine-grained appliance-level energy consumption in smart homes and building applications. However, the increasing adoption of behind-the-meter (BTM) energy sources such as solar panels and battery storage poses new challenges for conventional NILM methods that rely solely on at-the-meter data. The energy injected from the BTM sources can obscure the power signatures of individual appliances, leading to a significant decrease in NILM performance. To address this challenge, we present DualNILM, a deep multi-task learning framework designed for the dual tasks of appliance state recognition and injected energy identification. Using a Transformer-based architecture that integrates sequence-to-point and sequence-to-sequence strategies, DualNILM effectively captures multiscale temporal dependencies in the aggregate power consumption patterns, allowing for accurate appliance state recognition and energy injection identification. Extensive evaluation on self-collected and synthesized datasets demonstrates that DualNILM maintains an excellent performance for dual tasks in NILM, much outperforming conventional methods. Our work underscores the framework’s potential for robust energy disaggregation in modern energy systems with renewable penetration. Synthetic photovoltaic augmented datasets with realistic injection simulation methodology are open-sourced at [link].
1. Introduction
Non-Intrusive Load Monitoring (NILM) technique enables disaggregation of total household energy consumption into individual appliance usage patterns (Hart, 1992), offering a cost-effective solution for fine-grained energy monitoring. This capability is increasingly critical for smart grid applications, facilitating targeted energy-saving recommendations, demand response programs, and enhanced grid efficiency (Zoha et al., 2012).
Conventional NILM methods operate on a key assumption: the aggregate power measurements from the utility meter represent the sum of purely consumptive loads. Then these methods leverage unique electrical signatures of appliances to infer their operating states (or power) within the total load (Makonin and Popowich, 2015; Batra et al., 2014; Rafiq et al., 2024). However, the rapid adoption of behind-the-meter (BTM) renewable energy systems, particularly solar photovoltaics (PV), fundamentally disrupts this premise. The injection of energy from BTM sources creates bidirectional power flows that can dominate the aggregate power profile (especially during peak generation periods) (Saleem et al., 2022; Rafiq et al., 2024), severely obscuring the signatures of individual appliances and challenging the foundation of traditional NILM.
@@ 1. Introduction
The technical complexity of this challenge is significant. When PV generation exceeds instantaneous consumption, the disaggregation problem undergoes a fundamental mathematical transformation. Traditional NILM benefits from non-negativity constraints that bound the solution space. Each appliance contributes positively to the aggregate, creating a convex optimization landscape. In contrast, BTM generation shatters these constraints, expanding the feasible region to include arbitrary combinations of positive consumption and negative generation that yield identical net measurements. This transformation is further complicated by the electrical characteristics of modern grid-tied inverters, which operate at near-unity power factors to maximize generation efficiency. As a result, BTM injection almost purely perturbs the active power, whereas the measured reactive power is only weakly affected and remains largely governed by load behavior. This asymmetry renders conventional active-only NILM approaches particularly fragile under BTM injection.
Figure 1 empirically demonstrates these theoretical challenges through controlled laboratory experiments. Our micro-inverter injection setup reveals how even modest energy generation fundamentally distorts the aggregate signal. During injection periods, the expected additive structure of load aggregation breaks down. The aggregate signal (orange) deviates dramatically from the consumption-only baseline (blue), while the actual appliance states (bottom panels) become completely obscured. This controlled experiment confirms that the challenge transcends simple noise addition, revealing a fundamental incompatibility between conventional NILM architectures and the bidirectional power flows of modern energy systems.
From a utility and grid operator perspective, maintaining load visibility despite BTM resources is critical for demand response, grid stability, and planning. While smart inverters can theoretically report generation data, practical challenges include: ownership and privacy boundaries between customers and utilities, heterogeneous equipment with proprietary communication protocols, legacy installations lacking smart monitoring capabilities, and aggregated scenarios such as multi-tenant buildings or community solar arrangements without individual metering.
@@
Systematic characterization of BTM challenges: We analyze how behind-the-meter energy injection transforms the NILM problem from a constrained optimization to an under-determined inverse problem, formally identifying key failure modes such as signal eclipse and temporal aliasing.
@@
Comprehensive empirical validation: We demonstrate DualNILM’s state-of-the-art performance across diverse operational regimes with heterogeneous BTM injection sources, from controlled laboratory experiments (500W micro-inverter) to synthetic residential-scale scenarios (2kW PV systems).
Open research infrastructure: We release the synthetic PV-augmented versions (with realistic weather data) of REDD and UK-DALE datasets and the corresponding benchmark suites to support future research in this critical area [link]111https://github.com/MathAdventurer/PV-Augmented-NILM-Datasets.
2. Related work
@@
Non-Intrusive Load Monitoring (NILM) has been a subject of extensive research since its introduction by Hart (Hart, 1992). Traditional NILM approaches rely on identifying unique appliance signatures in the aggregate power signal to disaggregate individual appliance usage (Makonin and Popowich, 2015). Techniques such as Hidden Markov Models (HMMs) (Kim et al., 2011), factorial HMMs (Kolter and Jaakkola, 2012), and combinatorial optimization (Kolter and Johnson, 2011) have been widely used. With the advent of deep learning, neural network-based methods have significantly advanced NILM performance. Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) have been employed to capture temporal dependencies and extract features from power signals (Kelly and Knottenbelt, 2015a; Zhang et al., 2018; Ayub and El-Alfy, 2023). Sequence-to-point and sequence-to-sequence models have been proposed to improve disaggregation accuracy (Zhang et al., 2018; Du et al., 2016). Recently, researchers have also introduced advanced deep architectures, such as transformer-based models and graph neural networks, to further enhance disaggregation performance (Çavdar and Feryad, 2021; Shang and others, 2024). These approaches leverage long-range temporal context and inter-appliance relationships to achieve higher accuracy on benchmark datasets while improving the generalization ability of NILM models.
@@
NILM methods can also be categorized according to their task-oriented objectives. Many approaches focus on event-based detection, where the key aim is to recognize appliance switching events (ON/OFF transitions) and label those events with the corresponding device (Wang et al., 2022; Liaqat and Sajjad, 2022). This category typically relies on detecting changes in aggregate power that exceed a certain threshold, followed by a classification stage to identify which appliance caused the event. Various improvements to event detection have been proposed, including hybrid schemes that combine primary event triggers with secondary verification algorithms to reduce false detections (Lu and Li, 2019). By isolating transition points in time series, event-based NILM methods can be highly interpretable and computationally efficient, but may struggle with appliances that exhibit gradual power variations or multiple internal states (Meziane et al., 2017). On the other hand, the steady-state-based approaches emphasize continuous monitoring of power consumption, estimating the exact load of each device over time. These NILM methods often employ more complex models (e.g., HMMs or deep neural networks) and can capture multi-level operating states of appliances (Cao et al., 2023).
NILM Methods amid RE or BTM Integration
Previous work addressing NILM with the integration of renewable energy is limited. In (Hamidi and Gholizade-Narm, 2016), the authors investigate methods for managing power injection systems in renewable setups, emphasizing the stability of the system. Similarly, Aygen et al. (Aygen and Inci, 2020) consider a zero-sequence current injection technique to mitigate unbalanced loads in systems integrated with renewables, which, however, is not directly applicable to NILM. Only a few studies have attempted to explicitly disaggregate solar generation from the net metered signal. SunDance (Chen and Irwin, 2017) from Chen & Irwin presents a black-box approach specifically for solar disaggregation, combining physical solar generation models with machine learning to map weather conditions to solar output without appliance-level monitoring. Noori et al. (RA et al., 2023) recently examined solar disaggregation and highlighted that incorporating an external context, such as solar irradiance, can improve the detection of solar generation. In (Chen and Ardakanian, 2020), the authors investigate solar energy disaggregation but rely on high-frequency data and additional features, which may not be practical in real-world deployments. Li et al. (Li et al., 2024) propose BERT-based architectures for behind-the-meter PV monitoring, utilizing user embeddings and weather features for power regression tasks. Other studies have considered the use of additional sensors or prior knowledge of the energy injection (Abir et al., 2021), which limits their applicability.
@@
3. Problem Definition
Traditional NILM approaches infer appliance-level power consumption from aggregated household measurements under the assumption that all loads are strictly non-negative (Hart, 1992; Kolter and Johnson, 2011). However, modern residences increasingly incorporate BTM renewable generation, where micro-inverters inject power back to the grid, fundamentally violating this core assumption.
3.1. NILM with BTM Energy Injection
@@
where aggregates remaining loads and noise. While solving for is NP-hard (Hart, 1992), the non-negative structure historically enables acceptable performance through heuristic and learning-based methods (Makonin and Popowich, 2015; Batra et al., 2014).
@@
3.2. DNN-based NILM with Dual Tasks
Deep neural networks (DNNs) have shown promise in NILM by exploiting additional features or electrical signatures beyond active power (Kelly and Knottenbelt, 2015a; Zhang et al., 2018). Let denote the multi-dimensional feature vector at time , including active power, reactive power, current harmonics, etc. A neural model processes these features as:
@@ 3.2. DNN-based NILM with Dual Tasks
This formulation enables the model to learn complementary representations: discrete state transitions for appliances and continuous power variations for injection. By explicitly modeling injection rather than treating it as negative consumption, the framework naturally accommodates bidirectional power flows while reducing optimization complexity. The dual-task NILM formulation is also well-suited for DNN-based NILM models that can leverage rich features, capture complex patterns and dependencies in the data, allowing accurate estimation of the appliance states and the energy injection (Kelly and Knottenbelt, 2015a; Zhang et al., 2018; Kaselimi et al., 2019).
@@ 4.2. Model Architecture
where is the -th feature channel and is the corresponding encoded representation. Using separate CNN Encoders for each feature allows the model to learn feature-specific local patterns, enhancing its ability to capture the unique characteristics of different electrical measurements (Kelly and Knottenbelt, 2015a).
@@
The outputs of the CNN Encoders are concatenated along the feature dimension and fed into Transformer Encoders (Vaswani et al., 2017). The Transformer Encoder employs multi-head attention mechanisms to capture long-range dependencies in the input sequence:
where is the output of the Transformer Encoder. This component allows DualNILM to model complex interactions between different appliances and the energy injection source over time, which is crucial for accurate disaggregation and injection estimation (Zhang et al., 2018).
@@
The proposed architecture addresses the signal eclipse phenomenon directly through the above coupled design choices, ensuring both theoretical soundness and practical viability. First, the CNN-Transformer hybrid accommodates distinct temporal characteristics: appliance switching events exhibit sharp transients captured by convolutional kernels, while energy injection exhibits complex multiscale temporal dependencies and cross-correlation patterns requiring long-range attention mechanisms to capture their dynamics. Second, explicit separation of injection estimation through dedicated Seq2Seq projection prevents mathematical contamination of appliance state classification, where negative flows would corrupt binary decision boundaries. This architectural separation, combined with LayerNorm stabilization (Ba et al., 2016), enables direct processing of raw meter readings while maintaining modular scalability for incremental appliance addition. Third, the shared encoder representations enable cross-task synergy that significantly reduces the problem of identification, allowing accurate disambiguation between actual load reduction and injection-induced power dips within their respective constraint domains.
4.3. Training Objective and Optimization
@@
Our model can be optimized using gradient backpropagation approaches such as SGD (Amari, 1993), Adam (Kingma and Ba, 2014), etc. Specifically, we employ Dice loss (Milletari et al., 2016), which is particularly effective for imbalanced binary classification problems for appliance state recognition:
For injection estimation loss, we use the standard L2 measure: Root Mean Square Error (RMSE). Note that we can apply constraints/normalization on injection disaggregation targets as the maximum output of micro-inverters or community solar installations typically must adhere to standardized values like EN50583 (43).
5. Experimental Setup
@@ 5.1. Datasets and Simulation
We collected a real-world dataset in our laboratory and conducted extensive experiments. The dataset includes individual power consumption data for multiple appliances and the BTM energy injection from a micro-inverter. The data was collected from December 6 to December 15, 2023, and was sampled at 2-second intervals using sub-meters for each appliance on both active power and reactive power. The dataset includes the following appliances: i) Air purifier, ii) Heater, iii) Light bulb Type I, iv) Light bulb Type II, v) Air compressor, vi) Air conditioner, vii) Micro-inverter (energy injection). The laboratory data collection employed a combination of automated random control and manual random triggering to eliminate temporal correlations. For each of the included appliances, operating states were randomly selected, along with randomized run durations and random micro-inverter power level adjustments. This creates an adversarial scenario testing pure disaggregation capability without exploitable temporal patterns. Table 1 presents the key statistics of our laboratory dataset.
@@
To evaluate the generalizability of DualNILM and facilitate comparison with existing methods, we simulate energy injection scenarios using two widely-used public NILM datasets: REDD (Kolter and Johnson, 2011) and UK-DALE (Kelly and Knottenbelt, 2015b). Our simulation takes advantage of real-world solar irradiance data and photovoltaic (PV) system modeling to create realistic scenarios of residential solar energy injection. Detailed implementation is open-source and available in [link]. The simulation process involved the following steps:
@@
Acquisition of Solar Irradiance Data: We obtain historical solar irradiance data from the National Solar Radiation Database (NSRDB) provided by the National Renewable Energy Laboratory (NREL) (National Renewable Energy Laboratory (NREL), 2020). The NSRDB offers detailed solar radiation data, including Global Horizontal Irradiance (GHI), Direct Normal Irradiance (DNI), and Diffuse Horizontal Irradiance (DHI), with a temporal resolution of 30 minutes. We collect data for the rough geographical locations corresponding to the datasets:
@@
Simulation of Photovoltaic System Output: Using the collected irradiance data, we simulate a residential PV system with a designed maximum output capacity of 120 W, representing a small-scale household solar installation. The simulation accounts for the effects of temperature on the efficiency of PV, following the methods described in (Duffie and Beckman, 2013; Marion et al., 2002).
@@
(16) where is the rated capacity (120 W in our simulation) and is the inverter efficiency (0.96 based on (47)).
Reactive Power Simulation: For comprehensive power analysis, we simulate reactive power characteristics for both appliances and the PV system based on typical power factors from industry standards (22). For each appliance with active power , its reactive power is calculated as: